# FocusLLM: Scaling LLM’s Context by Parallel Decoding

### v1

@@ Abstract
Empowering LLMs with the ability to utilize useful information from a long context is crucial for many downstream applications. However, achieving long context lengths with the conventional transformer architecture requires substantial training and inference resources. In this paper, we present FocusLLM, a framework designed to extend the context length of any decoder-only LLM, enabling the model to focus on relevant information from very long sequences. FocusLLM processes long text inputs by dividing them into chunks based on the model’s original context length to alleviate the issue of attention distraction. Then, it appends the local context to each chunk as a prompt to extract essential information from each chunk based on a novel parallel decoding mechanism, and ultimately integrates the extracted information into the local context. FocusLLM stands out for great training efficiency and versatility: trained with an 8K input length with much less training cost than previous methods, FocusLLM exhibits superior performance across downstream long-context tasks and maintains strong language modeling ability when handling extensive long texts, even up to 400K tokens. Our code is available at https://github.com/leezythu/FocusLLM.
FocusLLM: Scaling LLM’s Context by Parallel Decoding
Zhenyu Li1, Yike Zhang1, Tengyu Pan1, Yutao Sun1, Zhichao Duan1, Junjie Fang2, Rong Han1, Zixuan Wang1, Jianyong Wang1thanks: Corresponding author 1Tsinghua University 2Xiamen University
@@ 1 Introduction
The importance of extending the context length of large language models (LLMs) cannot be overstated. In numerous applications, ranging from complex document analysis to generating coherent long-form text, the ability to effectively utilize extended context is critical. For instance, in tasks such as document summarization and question answering over lengthy articles, a more extensive context allows for a more comprehensive understanding and accurate responses Li et al. (2024a). However, leveraging long contexts in LLMs presents several formidable challenges. (1) The computational complexity of transformers Vaswani et al. (2017) grows quadratically with the sequence length, rendering the training process prohibitively expensive. (2) LLMs exhibit poor extrapolation performance for longer sequences, even after additional fine-tuning Chen et al. (2023a); Peng et al. (2023). (3) Acquiring high-quality long-text datasets, which are essential for training and fine-tuning, is exceedingly difficult Xiong et al. (2023); Wang et al. (2022).
@@
Figure 1: A comparison between FocusLLM and previous context scaling methods on the passkey retrieval task, including CEPE Yen et al. (2024), LongLLaMA Tworkowski et al. (2024) and Activation Beacon Zhang et al. (2024a). Our method extrapolates beyond the original context length of LLaMA, achieving 99% accuracy at a context length of 400K, with less training cost.
To circumvent the substantial costs of directly scaling the window length by fine-tuning on longer inputs, many approaches have attempted to modify the attention mechanism Xiao et al. (2023); Han et al. (2023) or compress tokens Zhang et al. (2024a); Chevalier et al. (2023); Ge et al. (2023) to theoretically achieve infinite length. While these methods can maintain lower perplexity over extended texts, the loss of information from earlier parts of the text hampers the model’s ability to perform precise understanding tasks such as information verification or question answering Zhang et al. (2024a).
Figure 2: The standard autoregressive model architecture. Suppose the input consists of a lengthy segment of information (e.g., a long document) and a user instruction (e.g., a question), with the total length being $L$. Due to the attention mechanism, the computational complexity is $O(L^{2})$, and $L$ is constrained by the length to which the model has been pre-trained. The performance on texts exceeding the length $L$ tends to deteriorate sharply.
@@
Figure 3: The FocusLLM framework. A small fragment of the local context is appended to each chunk (where $k$ = 3), which consists of memory tokens. The representations of the candidate tokens, obtained through parallel decoding, are then concatenated and integrated as memory back into the local context.
In this paper, we propose a training efficient and effective solution that enables large language models (LLMs) to focus on relevant content within long-context documents and generate high-quality outputs, called FocusLLM, by utilizing parallel decoding. FocusLLM is designed to adapt to any decoder-only language model with the modifications: i) the original model parameters are frozen to retain their generalization capabilities; ii) A small number of trainable parameters are added to enable the model to aggregate information from parallel decoding of different chunks. Not only does FocusLLM maintain low perplexity on long texts, but it also performs well on downstream tasks requiring precise understanding. Our evaluation on a widely adopted passkey retrieval task (Figure 1) has shown that FocusLLM, in contrast to the recently proposed context scaling methods, can sustain the accuracy of text comprehension on much longer texts.
FocusLLM bears several standout features: (1) Length Scaling: FocusLLM breaks the inherent positional limitations, allowing the model to handle text lengths that are expanded by tens or even hundreds of times. (2) Training Efficiency: Unlike full fine-tuning, we keep the original model parameters frozen and add only a small number of trainable parameters. The training process can be completed within a 0.5B tokens training budget, which is significantly smaller than that of previous works. (3) Versatility: FocusLLM not only excels in downstream tasks with clear instructions, such as question answering, but also demonstrates strong language modeling capabilities over long documents.
@@
We employ the FocusLLM framework to LLaMA-2-7B Touvron et al. (2023b), which has a default context length of 4K. To validate the effectiveness of FocusLLM, we evaluate it across a variety of tasks. Initially, we assessed FocusLLM’s language modeling capability. Trained with only 8K input length, FocusLLM maintains low perplexity on documents comprising 128K tokens and even longer sequences. Subsequently, to comprehensively evaluate the applicability of FocusLLM in real-world scenarios, we utilized two widely used benchmarks: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Longbench encompasses a diverse range of tasks, while $\infty$-Bench requires models to perform inference on extremely long sequences (>100K tokens). Experimental results demonstrate that FocusLLM has achieved superior performance on both benchmarks, surpassing all baselines including length extrapolation models, continual training models, and similar models designed for extreme long sequences.
In conclusion, FocusLLM is a training-efficient framework capable of achieving effective understanding and reasoning on long sequences at a minimal cost. We hope that FocusLLM can contribute to research related to developing easy-to-implement long-context capabilities.
2 Methodology
In this section, we introduce the design methodology of FocusLLM. First, we describe how we construct FocusLLM based on the architecture of LLM to enable it to handle extremely long text contexts. Then, we explain the training process of FocusLLM.
2.1 Architecture
As illustrated in Figure 2, the standard model architecture has a quadratic complexity and a corresponding limited context length. This limitation restricts the model’s application to longer texts, and FocusLLM is designed to address these drawbacks.
@@ 2.1 Architecture
The overall framework of FocusLLM is simple and intuitive. Each decoder in Figure 3 shares the same model (e.g. LLaMA-2). Besides, for the decoder handling each chunk, we augment the original decoder with a small set of additional parameters.
2.1.1 Notations.
@@
Given a long sequence with $S$ tokens $\{x_{1},...,x_{S}\}$, we segment them into memory tokens $\{x_{1},...,x_{m}\}$ and local tokens $\{x_{m+1},...,x_{S}\}$, with the length of local tokens not exceeding the model’s default context length, denoted as $L$. Concurrently, we divide the memory into chunks, labeled as $C_{1},C_{2},...,C_{k}$, with each chunk’s size also not exceeding $L$. These chunks can represent distinct documents or constitute a single long document.
We define the original decoder model as $F_{dec}$ and its hidden dimension $d_{dec}$. When processing the memory, to endow the model with the capability to generate candidate tokens, we introduce a small set of new parameters, resulting in the modified model $F^{\prime}_{dec}$. The candidate token is denoted as the trainable hidden states corresponding to the last local token $x_{S}$ in each chunk, serving as a signal indicating whether this chunk contains information relevant to the local context, as well as facilitating the prediction of the next token based on the current chunk.
2.1.2 Local Context Injection.
The aim of our FocusLLM is to distribute the burden of understanding long texts across each chunk. Therefore, unlike previous approaches that encode memory tokens and save them as cache, we append a small fragment of local tokens (we refer to it as the ‘prompt’ in the figure) behind each chunk and perform parallel decoding within each chunk, as shown in Figure 3. This strategy allows for a more efficient handling of long sequences by focusing the computational effort on relevant segments of the text, while maintaining the global information necessary for accurate decoding. We can formally define this process as follows:
@@
Here $j$ is a hyperparameter that determines the number of local tokens appended to each chunk. To reduce computational overhead, there is no need to append all local tokens. For instance, if the local context length is 2K, we could concatenate only 512 tokens or just a short instruction. In experiments, we adopt a default length of 512 tokens for inference, which is sufficient to encapsulate the necessary local contextual information.
2.1.3 Parallel Decoding.
FocusLLM transforms the next token prediction process based on long sequences into the simultaneous generation of candidate tokens from different chunks, followed by aggregating these candidate tokens as memory to produce the final token. We refer to this process as parallel decoding.
Differences from previous encoding methods. Here, we will elucidate the meaning of the term ‘decoding’ and explain the logical differences from previous encoding methods. Many previous methods involve encoding the context into some form of memory and retaining the key-value (KV) cache for subsequent use Yen et al. (2024); Tworkowski et al. (2024); Zhang et al. (2024a). For instance, CEPE Yen et al. (2024) divides long texts into chunks and then encodes each chunk with a smaller language model. FocusLLM also divides long texts into chunks, but within each chunk, FocusLLM undergoes a decoding process.
Specifically, the purpose of encoding is to transform a segment of text into hidden states for later use, whereas decoding is about generating the next token based on the current text. From this perspective, our method can be viewed as a form of implicit decoding. This is because our objective is not to acquire representations for entire chunks of text but rather to maintain information about the next token (based on the current chunk $\hat{C_{i}}$) via a special candidate token. Although we do not explicitly map the hidden states of the special candidate tokens through a softmax layer to the corresponding tokens in the vocabulary, our approach is fundamentally aligned with the concept of decoding. For further discussion, please refer to Appendix A.
How to obtain the representations of candidate tokens. Inspired by Zhang et al. (2024a), in order to preserve the generalizability of the original model as much as possible, we only add a new set of trainable parameters to the linear projection matrices of each layer:
$\{W^{\prime}_{Q},W^{\prime}_{K},W^{\prime}_{V},W^{\prime}_{O}\}_{l}$ (2)
where $W^{\prime}_{Q}$, $W^{\prime}_{K}$, $W^{\prime}_{V}$, and $W^{\prime}_{O}$ represent the new parameters for the query, key, value, and output matrices, respectively, and $l$ denotes the layer number. The modified decoder model are denoted as $F^{\prime}_{dec}$.
We apply parallel decoding on each chunk and obtain one candidate token from every chunk:
$e_{i}=F^{\prime}_{dec}(\hat{C_{i}})$ (3)
where $e_{i}$ consists of key-value hidden states $K_{e}$ and $V_{e}$ of the last token in each layer. More formally, the output of the candidate token in the self-attention module is calculated similarly to Zhang et al. (2024a):
$Q_{e}\leftarrow HW^{\prime}_{Q}\quad K_{e}\leftarrow HW^{\prime}_{K}\quad V_{e}\leftarrow HW^{\prime}_{V}\quad$ (4)
$A\leftarrow softmax\left(Q_{e}\left(K\oplus K_{e}\right)^{T}\right)$ (5)
$O_{e}\leftarrow V_{e}{W^{\prime}_{O}}^{T}\quad V_{e}\leftarrow A\left(V\oplus V_{e}^{T}\right)$ (6)
where $H\in R^{d_{dec}}$ is the input hidden states of the candidate token, $\oplus$ represents the concatenation of matrices, and $K,V$ correspond to the representations of the normal tokens in one chunk. Finally, the generated candidate tokens are concatenated with the local tokens and are subsequently processed by a frozen decoder. To optimize the candidate tokens, we design two loss functions of language modeling in Section 2.3. Note that the process of obtaining the candidate token from each chunk is independent, which allows us to perform parallel forwarding for each chunk. This is the reason we refer to this process as ‘parallel decoding’.
2.1.4 Efficiency.
2.2 Data
@@
To ensure the generalizability of our method and to maintain fairness in comparison with the baselines, we leverage the same training corpus as Activation Beacon Zhang et al. (2024a), with sequence lengths varying between 3K and 8K tokens. Specifically, we utilized 70K samples from the pre-training dataset RedPajama Together (2023b) and 10K samples from the instruction-following dataset LongAlpaca Chen et al. (2023b). RedPajama is an open-source pre-training dataset for LLaMA-1 Touvron et al. (2023a), while LongAlpaca is an instruction-following dataset for question answering. Both datasets are widely utilized in previous work. The detailed statistics of our training data are reported in Table 1.
Table 1: Length distribution of training corpus.
Length 2Ksimilar-$\sim$4K 4Ksimilar-$\sim$6K 6Ksimilar-$\sim$8K Total
Count 30K 16K 34K 80K
Portion 38% 20% 42% 100%
2.3 Training
We primarily conduct experiments on the LLaMA2-7B-Chat model. After incorporating the parameters mentioned in Section 2.1, the additional parameters amount to only 2B approximately. We train the model using an auto-regressive approach:
@@ 2.3 Training
Auto-Regressive Loss. FocusLLM is trained using a natural auto-regressive method. Specifically, we train the model to predict the next token, which encourages the candidate token to aggregate useful information from each chunk. Therefore, the loss is only applied to the local tokens:
$\min_{F^{\prime}_{dec}}-\sum_{i=2}^{L}\log(p(x_{i}\mid e_{1},\ldots,e_{k},x_{1},\ldots,x_{i-1};F^{\prime}_{dec}))$ (7)
In Section 2.1.1, we noted that local tokens are positioned subsequent to the memory tokens. However, to explicitly train the model to harness its ability to recover information from the memory tokens, we also set the local tokens to be a substring of the memory tokens. Specifically, based on the different selection methods for local tokens, we design two types of loss functions for joint training. i) Firstly, if the last $L$ tokens from a long document are selected as local tokens, with the remainder serving as memory tokens, we term this loss the ‘Continuation loss’, as it trains the model’s ability to naturally continue generating new tokens based on the context. ii) Alternatively, if we take the entire long document as memory and then randomly select $L$ continuous tokens from it as local tokens, we define this loss the ‘Repetition loss’, because it trains the model’s ability to repeat when clear information from the context is already available. Subsequent experiments have proven that both types of loss are important.
Generalizing Chunk Size. To ensure the model exhibits robust generalizability across various chunk sizes and number of candidate tokens, we maintain a constant local context size of 2048, while the chunk size is randomly selected from the set {64, 128, 256, 1024, 2048} during training.
Table 2: Language Modeling Assessment: perplexity analysis of various context scaling methods on the PG19, Proof-Pile, and CodeParrot. FocusLLM successfully extends context of the original Llama model and maintains low perplexity on extremely long sequences.

### v2

# FocusLLM: Precise Understanding of Long Context by Dynamic Condensing

@@ Abstract
Empowering LLMs with the ability to precisely understand long contexts is crucial for many downstream applications. However, handling long contexts with conventional transformer architecture requires substantial training and inference resources. Existing context condensing methods cannot accurately understand the full context, as there is a considerable amount of information loss in the condensing process. To address these issues, we present FocusLLM, a framework designed to extend the fixed context length of any decoder-only LLM, allowing the model to focus on relevant information from very long sequences. FocusLLM first divides long text input into chunks based on the model’s original context length. It then employs the dynamic condensing process to distill crucial information from each chunk. Ultimately, through the novel parallel decoding mechanism, FocusLLM can integrate the extracted information into its local context. FocusLLM stands out for great training efficiency and versatility: trained with an 8K input length and with much less training cost than previous methods, FocusLLM exhibits superior performance across downstream tasks and maintains strong language modeling ability when handling extensive long texts, even up to 400K tokens. Our code is available at https://github.com/leezythu/focusllm.
FocusLLM: Precise Understanding of Long Context by
Dynamic Condensing
Zhenyu Li, Yike Zhang, Tengyu Pan, Yutao Sun, Zhichao Duan, Junjie Fang, Rong Han, Zixuan Wang, Jianyong Wangthanks: Corresponding author Tsinghua University
@@ 1 Introduction
The importance of extending the context length of large language models (LLMs) cannot be overstated. In numerous applications, ranging from complex document analysis to generating coherent long-form text, the ability to effectively utilize extended context is critical. For instance, in tasks such as document summarization and question answering over lengthy articles, a more extensive context allows for a more comprehensive understanding and accurate responses. However, leveraging long contexts in LLMs presents several formidable challenges. (1) The computational complexity of transformers Vaswani et al. (2017) grows quadratically with the sequence length, rendering the training process prohibitively expensive. (2) LLMs exhibit poor extrapolation performance for longer sequences, even after additional fine-tuning Chen et al. (2023a); Peng et al. (2023). (3) Acquiring high-quality long-text datasets, which are essential for training and fine-tuning, is exceedingly difficult Xiong et al. (2023); Wang et al. (2022).
@@
Figure 1: A comparison between FocusLLM and previous context scaling methods on the passkey retrieval task, including CEPE, LongLLaMA and Activation Beacon. Our method extrapolates beyond the original context length of LLaMA, achieving 99% accuracy at a context length of 400K, with less training cost.
To circumvent the substantial costs of directly scaling the window length by continual training on longer inputs, recent work has proposed to drop unimportant tokens and retain important tokens, either by modifying the attention mechanism Xiao et al. (2023); Han et al. (2023) or by compressing the context into some specialized tokens Zhang et al. (2024a); Chevalier et al. (2023); Ge et al. (2023), in order to effectively condense long textual information. However, these methods overlook the fact that token importance changes dynamically during the decoding process: tokens previously considered unimportant may become crucial in later decoding steps. As a result, they share a common drawback, which we refer to as information loss: some tokens that will be needed in the future have already been discarded. For example, in Passkey Retrieval task Mohtashami and Jaggi (2024) illustrated in Figure 1, as the context length increases, the compression method Activation Beacon fails to retrieve passkey pairs that appeared in the earlier context.
@@
Figure 2: One decoding step of the FocusLLM framework. A small fragment of the local context (denoted as the dynamic prompt) is appended to each chunk. The representations of the candidate tokens, obtained through dynamic condensing and parallel decoding, are then concatenated and integrated back into the local context.
Considering the above issues, the question arises: can we extend the context length of an existing LLM at a low cost without any information loss? In this paper, we propose a training efficient and effective solution FocusLLM, which can maintain a precise understanding of the whole long context. Specifically, FocusLLM first divides a long text into chunks based on the model’s original context length. Then, the dynamic condensing process is applied, which appends dynamic prompts to each chunk to extract crucial information, ensuring no information loss. Finally, we use parallel decoding mechanism to aggregate information from different chunks and generate the next token. The original model parameters are kept frozen to maintain generalization capabilities, with only a small number of trainable parameters introduced for dynamic condensing.
@@
We employ the FocusLLM framework to the widely used LLaMA-2-7B model Touvron et al. (2023b), which has a default context length of 4K. In terms of efficiency, FocusLLM is trained on sequences shorter than 8K tokens and only requires a training budget of 0.5B tokens. To validate the effectiveness of FocusLLM, we evaluate it across a variety of tasks. Initially, we assessed FocusLLM’s language modeling capability. FocusLLM maintains low perplexity on documents comprising 128K tokens and even longer sequences. Subsequently, to comprehensively evaluate the applicability of FocusLLM in real-world scenarios, we utilized two widely used benchmarks: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Experimental results demonstrate that FocusLLM has achieved superior performance on both benchmarks, surpassing all baselines including length extrapolation models, continual training models, and similar models designed for extreme long sequences. The main contributions of this paper can be summarized as follows:
  • We propose the FocusLLM framework, which leverages novel dynamic condensing and parallel decoding mechanisms to avoid information loss and achieve precise understanding of long contexts, as shown in Figure 1.
  • Compared to previous context-scaling methods, FocusLLM achieves remarkable results with high training efficiency by introducing only a small set of trainable parameters and utilizing a training budget of 0.5B tokens.
  • Through comprehensive evaluation, FocusLLM outperforms all baselines on downstream tasks while maintaining low perplexity, demonstrating that it can seamlessly serve as a general-purpose language model.
2 Architecture
@@
The overall framework of FocusLLM is presented in Figure 2. Each decoder in the figure shares the same model (e.g. LLaMA-2).
2.1 Notations
@@
Given a long sequence with $S$ tokens $\{x_{1},...,x_{S}\}$, we segment them into memory tokens $\{x_{1},...,x_{m}\}$ and local context $\{x_{m+1},...,x_{S}\}$, with the length of local context not exceeding the model’s default context length, denoted as $L$. Concurrently, we divide the memory tokens into chunks, labeled as $C_{1},C_{2},...,C_{k}$, with each chunk’s size also not exceeding $L$. These chunks can represent distinct documents or a single long document. We define the original decoder model as $F_{\text{dec}}$ and its hidden dimension $d_{dec}$. To endow the model with the capability for dynamic condensing, we introduce a small set of new parameters, resulting in the modified model $F^{\prime}_{\text{dec}}$.
2.2 Dynamic Condensing
As highlighted in the introduction, the importance of tokens in the context dynamically changes at each decoding step. Previous work that condenses context using a fixed pattern suffers from the drawback of information loss. To address this issue, we propose the dynamic condensing mechanism, which consists of two key steps: dynamic prompt injection and candidate token generation.
Dynamic Prompt Injection. We append a small fragment of local context (we refer to it as the dynamic prompt in Figure 2) behind each chunk. The motivation is to aggregate the most critical information from each chunk for the current decoding step. We can formally define this process as follows:
@@ 2.2 Dynamic Condensing
Here $j$ is a hyperparameter that determines the number of local tokens appended to each chunk. We adopt a default length of 512 tokens for inference, which is sufficient to encapsulate the necessary local contextual information.
The last token of the dynamic prompt is used to generate candidate tokens, which we will explain in detail later. After each decoding step, when FocusLLM generates the next token, this token will be appended to the dynamic prompt  [^1] The first token of the dynamic prompt can be dropped to maintain its fixed length.. This updated dynamic prompt is then used to generate new candidate tokens in the next decoding step.
The dynamic prompt evolves with each decoding step, ensuring that the model always has access to the most relevant information for the current step.
Candidate Token Generation. Building on the dynamic prompt injection described above, we introduce candidate tokens to condense the information from each chunk that is crucial for the current decoding step. The candidate token is denoted as the trainable hidden states corresponding to the last local token $x_{S}$ in each chunk $\hat{C_{i}}$. To obtain the representations of candidate tokens, motivated by Zhang et al. (2024a), we add a new set of trainable parameters to the linear projection matrices of each layer, while keeping the original model parameters frozen to preserve its original decoding ability. Formally, the trainable parameters for dynamic condensing are:
$\{W^{c}_{Q},W^{c}_{K},W^{c}_{V},W^{c}_{O}\}_{l}$ (2)
where $W^{c}_{Q}$, $W^{c}_{K}$, $W^{c}_{V}$, and $W^{c}_{O}$ represent the new linear projections for the query, key, value, and output matrices associated with the candidate token, and $l$ denotes the layer number. The output of the candidate token in the self-attention module can be calculated as:
$Q_{c}\leftarrow H_{c}W^{c}_{Q}\quad K_{c}\leftarrow H_{c}W^{c}_{K}\quad V_{c}\leftarrow H_{c}W^{c}_{V}\quad$ (3)
$A_{c}\leftarrow\text{softmax}\left(Q_{c}\left(K\oplus K_{c}\right)^{T}\right)$ (4)
$O_{c}\leftarrow V_{c}{W^{c}_{O}}^{T}\quad V_{c}\leftarrow A_{c}\left(V\oplus V_{c}\right)^{T}$ (5)
where $H_{c}\in\mathbb{R}^{d_{dec}}$ is the input hidden state of the candidate token, $\oplus$ represents the concatenation of matrices, and $K,V$ correspond to the representations of the normal tokens in one chunk.
2.3 Parallel Decoding
Through the dynamic condensing process described above, we obtain one candidate token for each chunk. Notably, the process of obtaining the candidate token from each chunk is independent, enabling parallel forwarding for all chunks. Then the key/value representations of the candidate tokens are concatenated with the tokens in the local context layer by layer, as shown in Figure 2, and are finally processed by a frozen decoder to generate the next token.
We formally define the process of simultaneously generating candidate tokens from different chunks and then aggregating these candidate tokens to produce the final token as parallel decoding. This mechanism not only enables precise understanding of long contexts but also reduces the Transformer’s original $O(L^{2})$ computational complexity to $O((L/n)^{2})$. A detailed efficiency analysis is provided in Appendix A.
3 Training
@@
Regarding training data, to ensure the generalizability of our method and maintain fairness in comparison with the baselines, we leverage RedPajama Together (2023b) as the training corpus and sample examples with sequence lengths varying between 3K and 8K tokens from it. RedPajama is an open-source pre-training dataset for LLaMA-1 Touvron et al. (2023a), which is widely utilized in previous work Zhang et al. (2024a); Yen et al. (2024). Detailed statistics are reported in Appendix B.
@@
Auto-Regressive Loss. Specifically, we train the model to predict the next token, and the loss is only applied to tokens in the local context, which encourages the candidate token to aggregate useful information from each chunk.
$\min_{F^{\prime}_{\text{dec}}}-\sum_{i=2}^{S-m}\log(p(x_{m+i}\mid c_{1},\ldots,c_{k},x_{m+1},\ldots,x_{m+i-1}))$ (6)
Here, $c_{i}$ represents the candidate token generated by the $i$-th chunk. Specifically, based on the relationship between the memory tokens $\{x_{1},...,x_{m}\}$ and the local context $\{x_{m+1},...,x_{S}\}$, we design two loss functions for joint training. i) If the local context is a continuation of the memory tokens, we term this loss the Continuation Loss, as it trains the model to naturally generate new tokens that follow the given context. ii) Alternatively, if we randomly select $L$ consecutive memory tokens as local context, we define this loss as the Reconstruction Loss, as it trains the model to reconstruct tokens when clear contextual information is available. Subsequent experiments demonstrate that both types of loss are essential.
@@
Table 1: Language Modeling Assessment: perplexity analysis of various context scaling methods on the PG19, Proof-Pile, and CodeParrot. FocusLLM successfully maintains low perplexity on extremely long sequences.